OpenStudy (anonymous):

How is F'' of (1-x)e^-x Equal to (x-2)e^-x ?

OpenStudy (anonymous):

if F(x) = (1-x)e^-x then the double differential is that, using F' = u(dv/dx) + v(du/dx) where u= (1-x) v=e^-x then do the same method again to work out the F''

OpenStudy (anonymous):

Ok product rule then, well its still not working!

OpenStudy (anonymous):

I get (0)(e^-x)+(1-x)(e^-x)

OpenStudy (anonymous):

u=(1-x) du/dx=-1 v=e^-x dv/dx=-e^-x

OpenStudy (anonymous):

omg of course :) thanks!

OpenStudy (anonymous):

but wait how come the - sign can be used 2x?

OpenStudy (anonymous):

i dont understand your q

OpenStudy (anonymous):

I have (1-x^1) I bring the 1 down in front of the x and its (1-1)?

OpenStudy (anonymous):

how does it become (1-(-1)?

OpenStudy (anonymous):

the differential of a constant is zero , 0

OpenStudy (anonymous):

ok so its -1 then not -2? where does the -2 come from in (x-2)?

OpenStudy (anonymous):

you have to differentiate again F' is dy/dx F'' is d^2y/dx^2 F''' is d^3y/dx^3 ;;; ;;; F^n is d^ny/dx^n

OpenStudy (anonymous):

yes I am diferentiating again F'' to me is I get (0)(e^-x)+(1-x)(e^-x) using the product rule?

OpenStudy (anonymous):

ok what about (0)(-1)e^-x*(1-x)-e^-x

OpenStudy (anonymous):

Can you please tell me what to do I can't spend 30 min on this one stupid derivative!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh the original question is xe^-x sad part is this is an example in the book I don't even need to do it I just want to understand it

OpenStudy (anonymous):

let u=1-x u'=-1 v=e^-x v'=-e^-x then uv'+vu'= [1-x](-)e^-x + (-1)e^-x so e^-x[(x-1)-1] e^-x [x-2] then again do the same thing let u=e^-x u'=-e^-x v=x-2 v'=1 uv'+vu' = e^-x[1] + [x-2](-1)e^-x = e^-x *[ 1+2-x] = e^-x *[3-x] mmm, can you spot where i went wrong?

OpenStudy (anonymous):

How does u=1-x equal u'=-1 I get 0 when I do that

OpenStudy (anonymous):

oh crap nevermind! UGH I hate this sometimes. the constant is factored out

OpenStudy (anonymous):

the differential of -x is -1*x so we are left with -1 if you choose u=-1 v=x then v'=1 u'=0 uv'+vu' = (-1)*1 + x*(0) = -1

OpenStudy (anonymous):

Last question, for f' of xe^-x how does the x and the -1 get grouped like (1-x) When I get f' I get (-1)(x)e^-x

OpenStudy (anonymous):

u=x u'=1 v=e^-x v'=-e^-x uv'+vu'= x(-1)e^-x + (1)*e^-x = e^-x* [-x+1]

OpenStudy (anonymous):

Thanks Rezz you are the man or woman!

OpenStudy (anonymous):

yw!! im a man btw PS, if you get stuck with differentiation always use product rule, it also works with using in fractions, so for example you dont need to use the quotient rule just product rule, im in 2nd uni doing maths and i stil dont use quotient rule!!! ^ | was for your own information!

OpenStudy (anonymous):

I didnt realize that when using u'(g)+u(g') you could use the chain rule on "g" first and THEN you get g'. I was splitting g' up into -x and then e^-x as two seperate things, if this makes any sense

OpenStudy (anonymous):

you can use quotient rule for the first question where e^-x = 1/e^x

OpenStudy (anonymous):

no idea!

OpenStudy (anonymous):

haha thats ok